OpenStudy (anonymous):

how do u solve 3 over 4y + 11 over 16 is greater than 19 over 16

OpenStudy (anonymous):

(3/4y)+(14/11)>(19/16)?

OpenStudy (anonymous):

m' I right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

I help some one I will be back

OpenStudy (anonymous):

3*11=33 4y*14=56y 4y*11=44y for denominator now we have [(33+56y)/44y]>(19/16) time and cross 17*(33+56y)>44y*19 528+896y > 896y-836y>-528 60y >-528 y >-528 /60 y >-44/5